Eddies age is 4 times macros. in 10 years eddie will be twice as old as marco will be. how old is each now
Answer by stanbon(75887)

Eddies age is 4 times macros. in 10 years eddie will be twice as old as marco will be. how old is each now?
----------
Equations:
E = 4M
(E+10) = 2(M+10)
---
Substitute for "E" and solve for "M":
(4M+10) = 2(M+10)
4M + 10 = 2M+20
2M = 10
M = 5 (Marco's age now)
Since E = 4M, E = 20 (Eddie's age now)
==========================================
Cheers,
San H.
================